OpenStudy (ilovebmth1234):

Which statement about the effusion rates of nitrogen and oxygen is true? A) Oxygen effusion 1.07 times faster than nitrogen B) nitrogen effusion 1.07 times faster than oxygen C) oxygen effusion 0.9 times faster than nitrogen D) nitrogen effusion 0.9 times faster than oxygen

OpenStudy (ilovebmth1234):

@Missiey

OpenStudy (anonymous):

:)

OpenStudy (ilovebmth1234):

:3

OpenStudy (whpalmer4):

the ratio of effusion rates is inversely proportional to the square root of the molar mass. \[\frac{D_1}{D_2} = \sqrt{\frac{M_2}{M_1}}\]